# FC-KAN: Function Combinations in Kolmogorov-Arnold Networks

### v2

@@ Abstract
In this paper, we introduce FC-KAN, a Kolmogorov-Arnold Network (KAN) that leverages combinations of popular mathematical functions such as B-splines, wavelets, and radial basis functions on low-dimensional data through element-wise operations. We explore several methods for combining the outputs of these functions, including sum, element-wise product, the addition of sum and element-wise product, representations of quadratic and cubic functions, concatenation, linear transformation of the concatenated output, and others. In our experiments, we compare FC-KAN with a multi-layer perceptron network (MLP) and other existing KANs, such as BSRBF-KAN, EfficientKAN, FastKAN, and FasterKAN, on the MNIST and Fashion-MNIST datasets. A variant of FC-KAN, which uses a combination of outputs from B-splines and Difference of Gaussians (DoG) in the form of a quadratic function, outperformed all other models on the average of 5 independent training runs. We expect that FC-KAN can leverage function combinations to design future KANs. Our repository is publicly available at: https://github.com/hoangthangta/FC_KAN.
@@ 1 Introduction
The work by Liu et al. [1, 2] helped to revive the usage of the Kolmogorov-Arnold representation theorem (KART) in neural networks. They highlighted the use of learnable activation functions as "edges" to fit training data better, in contrast to the traditional use of fixed activation functions as "nodes" in multi-layer perceptrons (MLPs). The foundation of KANs is based on KART, which was developed to address Hilbert’s 13th problem [3], specifically, the statement, “Prove that the equation of seventh degree $x^{7}+ax^{3}+bx^{2}+cx+1=0$ is not solvable by means of any continuous functions of only two variables”. KART asserts that any continuous function of multiple variables can be represented as a sum of continuous functions of a single variable [4]. Motivated by this idea, numerous researchers have worked on creating different types of KANs utilizing widely known polynomials and basis functions.
@@
While these works use a single function (not mentioning activation functions) in constructing KANs [5, 6, 7, 8], there are a few ones using function combinations [9, 10, 11, 2]. In several works that designed MLP-based neural networks, the function combinations typically occur in activation functions to improve model performance and stability [12, 13, 14]. In KANs, the combinations occur directly on the functions used to fit input data, rather than through activation functions. For instance, BSRBF-KAN [9], which combines B-splines and Gaussian Radial Basis Functions (GRBFs) in all network layers. However, the authors only applied addition operations to data tensors without incorporating element-wise multiplications, which we believe does not effectively capture data features. Additionally, element-wise multiplications may lead to memory issues or slow calculations when performed on high-dimensional data. While Liu et al. [2] used element-wise multiplications alongside additions in MultKAN to enhance data capture, their focus was limited to small-scale examples.
@@
With the purpose of exploiting data features efficiently, we propose a novel KAN, FC-KAN (Function Combinations in Kolmogorov-Arnold Networks), which leverages various functions to capture input data throughout the network layers and combines them in low-dimensional spaces, such as the output layer, using various methods mainly based on element-wise operations, including sum, product, the combination of sum and product, representations of quadratic and cubic functions, and concatenation. We avoid using a higher-degree function due to their increased computational demands on data tensors, which can lead to memory errors. As a result, the model is able to capture more data features, leading to improved performance on the MNIST and Fashion-MNIST datasets compared to other KAN networks. Additionally, employing n-degree functions aligns with the core concept of KAN, where they are used both to capture input data and to represent data features in the output. We expect this exciting development to drive the proliferation of basis function combinations in neural networks. In summary, our main contributions are:
@@
  • Construct FC-KAN to demonstrate the effectiveness of function combinations through various methods applied to low-dimensional data, such as model outputs.
  • Compare combination methods on the function outputs using mainly element-wise operations to search for proper methods.
@@
Aside from this section, the paper is organized as follows: Section 2 discusses related work on KART and KANs. Section 3 details our methodology, covering KART, the design of the KAN architecture, several existing KANs, and FC-KAN. Section 4 presents our experiments, comparing FC-KAN variants with MLP and other existing KANs using data from the MNIST and Fashion-MNIST datasets. Additionally, this section includes a comparison of combination methods within FC-KAN. We address some limitations of our study in Section 5. Lastly, Section 6 offers our conclusions and potential directions for future research.
@@ 2.1 KART and KAN
Despite the long history of KART’s application in neural networks, it had not garnered significant attention in the research community until the recent study by Liu et al. [1]. They suggested moving away from strict adherence to KART and generalizing it to develop KANs with additional neurons and layers. Our intuition aligns with this perspective as it helps to mitigate the issue of non-smooth functions when applying KART to neural networks. Consequently, KANs have the potential to outperform MLPs in both accuracy and interpretability for small-scale AI + Science tasks. However, the resurgence of KANs has also faced criticism from Dhiman [25], who argue that KANs are essentially MLPs with spline-based activation functions, in contrast to traditional MLPs with fixed activation functions.
By introducing a new perspective to the scientific community in the new design of neural networks, KANs inspired many works to prove their effectiveness by topics, including expensive problems [26], keyword spotting [27], mechanics problems [28], quantum computing [29, 30, 31], survival analysis [32], time series forecasting [33, 34, 35, 36, 37], and vision tasks [38, 39, 40]. Also, many novel KANs utilize well-known mathematical functions, particularly those capable of handling curves, such as B-Splines [41] (Original KAN [1], EfficientKAN [^1] https://github.com/Blealtan/efficient-kan, BSRBF-KAN [9]), Gaussian Radial Basis Functions (GRBFs) (FastKAN [5], DeepOKAN [28], BSRBF-KAN [9]), Reflection SWitch Activation Function (RSWAF) in FasterKAN [6], Chebyshev polynomials (TorchKAN [42], Chebyshev KAN [8]), Legendre polynomials (TorchKAN [42]), Fourier transform (FourierKAN [^2] https://github.com/GistNoesis/FourierKAN/, FourierKAN-GCF [43]), wavelets [7, 44], and other polynomial functions [45].
@@ 2.2 Function combinations in KANs and other neural networks
Several works utilize the function combinations to design novel KANs. Ta [9] mentioned the combination of functions – B-Splines and radial basis functions – in designing KANs. Their BSRBF-KAN showed better convergence on the training data for MNIST and Fashion-MNIST. Liu et al. [2] introduced MultKAN, which consists of multiplication operations to capture multiplicative structures in data better. Unlike KAN [1], which directly copies nodes, MultKAN uses both addition nodes (copied from subnodes) and multiplication nodes (multiplying multiple subnodes). However, their work focused on small-scale examples only. Yang et al. [10] utilized function combinations to create optimal activation functions at each node using an adaptive strategy, addressing the drawbacks of single activation functions in their S-KAN model. They also extended S-KAN to S-ConvKAN, which showed superior performance in image classification tasks, outperforming CNNs and KANs with comparable structures. In another work,  Altarabichi [11] proposed the replacement of the sum with the average function in KAN neurons that can improve the model performance and keep the training stability in their DropKAN [46].
@@
Before the appearance of KANs, several studies focused on constructing combinations of activation functions (along with their weights) or utilizing a set of flexible activations in neural networks to enhance model performance and stability. Jie et al. [12] introduced a new family of flexible activation functions for LSTM cells, along with another family developed by combining ReLUs and ELUs. Their findings demonstrate that LSTM models using the P-Sig-Ramp flexible activations significantly improve time series forecasting. Additionally, the P-E2-ReLU activation exhibits enhanced stability and performance in lossy image compression tasks with convolutional autoencoders. Xu and Chen [13] investigated a selection of widely used activation functions across various datasets in classification and regression tasks. They then created combinations of the best-performing activation functions within a convex restriction, showing improved performance compared to the corresponding base activation functions in a standard broad learning system. In another study, Zhang [14] developed a novel deep neural network (DNN) that optimizes activation function combinations for different neurons based on extensive simulations. The experiments showed that their DNNs outperformed those that rely on a single activation function. Instead of using the inner product between data vectors and weights,  Fan et al. [47] replaced this operation with a quadratic function in neurons within deep quadratic neural networks. Quadratic neurons provide enhanced expressive capability compared to conventional neurons, highlighting the advantages of quadratic networks in terms of expressive efficiency, unique representation, compact architecture, and computational capacity.
@@
In addition to exploring function combinations in mathematics, it is also possible to merge function outputs using tensor operations. When the data dimensions are the same, element-wise operations such as summation or product can be applied to combine these outputs [48] which appear in many multi-modal tasks [49, 50, 51]. However, the computational challenges posed by high-dimensional data, such as inefficiencies in tensor-product operations, have led to a growing body of work aimed at optimizing and accelerating these processes [52, 53]. Furthermore, other studies have focused on tensor fusion and tensor decomposition techniques to extract meaningful features from data tensors and to simplify data combinations [54, 55, 56, 57].
@@ 3.2 The design of KANs
Inspired by KART, Liu et al. [1] developed KANs but recommended extending the approach to incorporate greater network widths and depths. To address this, appropriate functions $\Phi_{q}$ and $\phi_{q,p}$ need to be identified. A typical KAN network with $L$ layers processes the input $\mathbf{x}$ to produce the output as follows:
@@ 3.3 Implementation of the current KANs
Figure 1: Left: The structure of KAN(2,3,1). Right: The simulation of how to calculate $\phi_{1,1,1}$ by control points and B-splines.
LiuKAN [^3] We refer to the original KAN as LiuKAN, following the first author’s last name [1], while another work [7] refers to it as Spl-KAN. was implemented by Liu et al. [1] by using the residual activation function $\phi(x)$ as the sum of the base function and the spline function with their corresponding weight matrices $w_{b}$ and $w_{s}$:
@@
where $b(x)$ equals to $silu(x)$ as in Equation 7 and $spline(x)$ is expressed as a linear combination of B-splines $B_{i}$s and control points (coefficients) $c_{i}$s as in Equation 8. Each activation function is activated with $w_{s}=1$ and $spline(x)\approx 0$, while $w_{b}$ is initialized by using Xavier initialization. Figure 1 shows the structure of KAN(2,3,1) and how to calculate an inner function $\phi$ by control points and B-splines.
@@
where $r=\|x-c\|$ is the distance between an input vector $x$ and a center $c$, and $\epsilon$ ($epsilon>0$) is a sharp parameter that controls the width of the Gaussian function. FastKAN uses a special form of RBFs, Gaussian RBFs where $\epsilon=\frac{1}{2}$ as [5]:
@@
where $w_{i}$ represents adjustable weights or coefficients, and $\phi$ denotes the radial basis function as in Equation 9.
FasterKAN outperforms FastKAN in both forward and backward processing speeds [6]. It uses Reflectional Switch Activation Functions (RSWAFs), which are variants of RBFs and straightforward to compute due to their uniform grid structure. The RSWAF function is shown as follows:
@@
Then, the RSWAF network with $N$ centers will be:
@@
where $b(x)$ and $w_{b}$ are the base function (linear) and its base matrix. $\phi_{BS}(x)$ and $\phi_{RBF}(x)$ are B-Spline and RBF functions, and $w_{s}$ is the spline matrix.
Wav-KAN is a neural network architecture that integrates wavelet functions into Kolmogorov-Arnold Networks to address challenges in interpretability, training speed, robustness, and computational efficiency found in MLP and LiuKAN [7]. By efficiently capturing both high and low-frequency components of input data, Wav-KAN achieves a balance between accurately representing the data structure and avoiding overfitting. The authors use several wavelet types, including the DoG, Mexican hat, Morlet, and Shannon. In our paper, we use the DoG function to combine other functions to create function combinations. The formula for DoG is:

### v3

@@ Abstract
In this paper, we introduce FC-KAN, a Kolmogorov-Arnold Network (KAN) that leverages combinations of popular mathematical functions such as B-splines, wavelets, and radial basis functions on low-dimensional data through element-wise operations. We explore several methods for combining the outputs of these functions, including sum, element-wise product, the addition of sum and element-wise product, representations of quadratic and cubic functions, concatenation, linear transformation of the concatenated output, and others. In our experiments, we compare FC-KAN with a multi-layer perceptron network (MLP) and other existing KANs, such as BSRBF-KAN, EfficientKAN, FastKAN, and FasterKAN, on the MNIST and Fashion-MNIST datasets. Two variants of FC-KAN, which use a combination of outputs from B-splines and Difference of Gaussians (DoG) and from B-splines and linear transformations in the form of a quadratic function, outperformed overall other models on the average of 5 independent training runs. We expect that FC-KAN can leverage function combinations to design future KANs. Our repository is publicly available at: https://github.com/hoangthangta/FC_KAN.
@@ 1 Introduction
Recently, the works by Liu et al. [1, 2] have gained significant attention from the research community for applying the Kolmogorov-Arnold representation theorem (KART) in neural networks through the introduction of KANs. They highlighted the use of learnable activation functions as "edges" to fit training data better, in contrast to the traditional use of fixed activation functions as "nodes" in multi-layer perceptrons (MLPs). The foundation of KANs is based on KART, which was developed to address Hilbert’s 13th problem [3], specifically, the statement, “Prove that the equation of seventh degree $x^{7}+ax^{3}+bx^{2}+cx+1=0$ is not solvable by means of any continuous functions of only two variables”. KART asserts that any continuous function of multiple variables can be represented as a sum of continuous functions of a single variable [4].
@@
Inspired by KANs [1, 2], numerous researchers have explored developing novel neural networks using widely known polynomials and basis functions. While most works use a single function in constructing KANs [5, 6, 7, 8], a few have investigated function combinations [9, 10, 11, 2]. In several works on MLP-based neural networks, function combinations typically appear in activation functions to enhance model performance and stability [12, 13, 14]. In KANs, the combinations occur directly on the basis functions used to fit input data, rather than through activation functions. For instance, Ta [9] designed BSRBF-KAN, which combines B-splines and Gaussian Radial Basis Functions (GRBFs) in all network layers. The authors only applied addition operations to data tensors without incorporating element-wise multiplications, which we believe does not effectively capture data features. Although Liu et al. [2] used element-wise multiplications alongside additions in MultKAN to enhance data capture, their focus was limited to small-scale examples. Another study developed ReLU-KAN by replacing B-splines with ReLU activations, along with addition and multiplication operations, to improve GPU parallelization and data fitting [15], but it focused solely on ReLU without exploring combinations with other functions.
@@
With the purpose of exploiting data features efficiently, we propose a novel KAN, FC-KAN (Function Combinations in Kolmogorov-Arnold Networks), which leverages various functions to capture input data throughout the network layers and combines them in low-dimensional spaces, such as the output layer, using various methods mainly based on element-wise operations, including sum, product, the combination of sum and product, representations of quadratic and cubic functions, and concatenation. We avoid using a higher-degree function due to their increased computational demands on data tensors, which can lead to memory errors. As a result, FC-KAN is able to capture more data features, leading to improved performance on the MNIST and Fashion-MNIST datasets compared to other KAN networks. Moreover, employing n-degree functions aligns with the core concept of KAN, where they are used both to capture input data and to represent data features in the output. We expect this exciting development to drive the proliferation of function combinations in neural networks, particularly in KANs. In summary, our main contributions are:
@@
  • Introduce FC-KAN, a novel KAN that explores function combinations through various methods applied to low-dimensional data.
  • Evaluate the performance of different combination methods in FC-KAN on two image classification datasets: MNIST and Fashion-MNIST.
  • Investigate whether model performance in full-data training can be inferred from models trained with limited data.
@@
Aside from this section, the paper is organized as follows: Section 2 discusses related work on KART and KANs. Section 3 details our methodology, covering KART, the design of the KAN architecture, several existing KANs, and FC-KAN. Section 4 presents our experiments, comparing FC-KAN variants with MLP and other existing KANs using data from the MNIST and Fashion-MNIST datasets. Besides, this section includes a comparison of combination methods within FC-KAN, a misclassification analysis, and an analysis of model performance with limited data. We mention some limitations of our study in Section 5. Lastly, Section 6 offers our conclusions and potential directions for future research.
@@ 2.1 KART and KAN
Despite the long history of KART’s application in neural networks, it had not garnered significant attention in the research community until the recent study by Liu et al. [1]. They suggested moving away from strict adherence to KART and generalizing it to develop KANs with additional neurons and layers. Our intuition aligns with this perspective as it helps to mitigate the issue of non-smooth functions when applying KART to neural networks. Consequently, KANs have the potential to outperform MLPs in both accuracy and interpretability for small-scale AI + Science tasks. However, KANs have also faced criticism from Dhiman [26], who argue that they are essentially MLPs with spline-based activation functions, in contrast to traditional MLPs with fixed activation functions. KANs also face the problem of using too many parameters compared to MLPs. Yu et al. [27] indicated that KANs are not better than MLPs when using the same number of parameters and FLOPs, except for symbolic formula representation tasks.
By introducing a new perspective to the scientific community in the neural network designs, KANs inspired many works to prove their effectiveness by topics, including expensive problems [28], keyword spotting [29], mechanics problems [30], quantum computing [31, 32, 33], survival analysis [34], time series forecasting [35, 36, 37, 38, 39], and vision tasks [40, 41, 42]. Also, many novel KANs utilize well-known mathematical functions, particularly those capable of handling curves, such as B-splines [43] (Original KAN [1], EfficientKAN [^1] https://github.com/Blealtan/efficient-kan, BSRBF-KAN [9]), Gaussian Radial Basis Functions (GRBFs) (FastKAN [5], DeepOKAN [30], and BSRBF-KAN [9]), Reflection SWitch Activation Function (RSWAF) in FasterKAN [6], Chebyshev polynomials (TorchKAN [44], Chebyshev KAN [8]), Legendre polynomials (TorchKAN [44]), Fourier transform (FourierKAN [^2] https://github.com/GistNoesis/FourierKAN/, FourierKAN-GCF [45]), wavelets [7, 46], and other polynomial functions [47].
@@ 2.2 Function Combinations in KANs and Other Neural networks
Several works utilize the function combinations to design novel KANs. Ta [9] mentioned the combination of functions – B-splines and radial basis functions – in designing KANs. Their BSRBF-KAN showed better convergence on the training data for MNIST and Fashion-MNIST. Liu et al. [2] introduced MultKAN, which consists of multiplication operations to capture multiplicative structures in data better. Unlike KAN [1], which directly copies nodes, MultKAN uses both addition nodes (copied from subnodes) and multiplication nodes (multiplying multiple subnodes). However, their work focused on small-scale examples only. Yang et al. [10] utilized function combinations to create optimal activation functions at each node using an adaptive strategy, addressing the drawbacks of single activation functions in their S-KAN model. They also extended S-KAN to S-ConvKAN, which showed superior performance in image classification tasks, outperforming CNNs and KANs with comparable structures. In another work,  Altarabichi [11] proposed the replacement of the sum with the average function in KAN neurons that can improve the model performance and keep the training stability in their DropKAN [48]. Unlike other studies, Qiu et al. [15] developed ReLU-KAN, replacing B-splines with a novel basis function that leverages matrix operations (addition and multiplication) and ReLU activations to enhance GPU parallelization and fitting performance.
@@
Before the appearance of KANs, several studies focused on constructing combinations of activation functions or utilizing a set of flexible activations in neural networks to enhance model performance and stability. Jie et al. [12] introduced a new family of flexible activation functions for LSTM cells, along with another family developed by combining ReLUs and ELUs. Their findings demonstrate that LSTM models using the P-Sig-Ramp flexible activations significantly improve time series forecasting. Additionally, the P-E2-ReLU activation exhibits enhanced stability and performance in lossy image compression tasks with convolutional autoencoders. Xu and Chen [13] investigated a selection of widely used activation functions across various datasets in classification and regression tasks. They then created combinations of the best-performing activation functions within a convex restriction, showing improved performance compared to the corresponding base activation functions in a standard broad learning system. In another study, Zhang [14] developed a novel deep neural network (DNN) that optimizes activation function combinations for different neurons based on extensive simulations. The experiments showed that their DNNs outperformed those that rely on a single activation function. Instead of using the inner product between data tensors and weights,  Fan et al. [49] replaced this operation with a quadratic function in neurons within deep quadratic neural networks. Quadratic neurons provide enhanced expressive capability compared to conventional neurons, highlighting the advantages of quadratic networks in terms of expressive efficiency, unique representation, compact architecture, and computational capacity.
@@
Some other works focus on combining function outputs using tensor operations, exploring various methods to aggregate outputs efficiently. When the data dimensions are the same, element-wise operations such as summation or product can be applied to combine these outputs [50], which appear in many multi-modal tasks [51, 52, 53]. However, the computational challenges posed by high-dimensional data, such as inefficiencies in tensor-product operations, have led to a growing body of work aimed at optimizing and accelerating these processes [54, 55]. Furthermore, other studies have focused on tensor fusion and tensor decomposition techniques to extract meaningful features from data tensors and to simplify data combinations [56, 57, 58, 59].
@@ 3.2 Design of KANs
Inspired by KART, Liu et al. [1] developed KANs but recommended extending the approach to incorporate greater network widths and depths. To address this, appropriate functions $\Phi_{q}$ and $\phi_{q,p}$ in Equation 1 need to be identified. A typical KAN network with $L$ layers processes the input $\mathbf{x}$ to produce the output as follows:
@@ 3.3 Implementation of the Current KANs
Figure 1: Left: The structure of KAN(2,3,1). Right: The simulation of how to calculate $\phi_{1,1,1}$ by control points and B-splines. $G$ and $k$ is the grid size and the spline order, the number of B-splines equals $G+k=3+3=6$.
Figure 2: The plots of the functions when fitted to pass through the 4 chosen points.
LiuKAN [^3] We refer to the original KAN as LiuKAN, following the first author’s last name [1], while another work [7] refers to it as Spl-KAN. was created by Liu et al. [1] by using the residual activation function $\phi(x)$ as the sum of the base function and the spline function with their corresponding weight matrices $w_{b}$ and $w_{s}$:
@@
where $b(x)$ equals to $silu(x)$ as in Equation 7 and $spline(x)$ is expressed as a linear combination of B-splines $B_{i}$s and control points (coefficients) $c_{i}$s as in Equation 8. Each activation function is activated with $w_{s}=1$ and $spline(x)\approx 0$, while $w_{b}$ is initialized by using Xavier initialization.
Figure 1 shows the architecture of KAN(2,3,1), which includes 2 input nodes, 3 hidden nodes, and 1 output node. The output of each node is derived from the sum of individual functions $\phi$, called "edges." The diagram also explains the computation of the inner function $\phi$ using control points (coefficients) and B-splines. The number of B-splines is determined by adding the grid size $G$ and the spline order $k$, resulting in $G+k=3+3=6$, corresponding to the range of $i$ from 0 to 5.
@@
where $r=\|x-c\|$ is the distance between an input vector $x$ and a center $c$, and $\epsilon$ ($epsilon>0$) is a sharp parameter that controls the width of the Gaussian function. FastKAN uses a special form of RBFs, Gaussian RBFs where $\epsilon=\frac{1}{2h^{2}}$ as [5]:
@@
FasterKAN outperforms FastKAN in both forward and backward processing speeds [6]. It uses Reflectional Switch Activation Functions (RSWAFs), which are variants of RBFs. RSWAFs are activation functions that are easy to compute because of their uniform grid structure. The RSWAF function is shown as follows:
@@
where $h$ controls the function width. Then, the RSWAF network with $N$ centers will be:
[…]
$=\sum_{i=1}^{N}w_{i}\left(1-\left(\tanh\left(\frac{||x-c_{i}||}{h}\right)\right)^{2}\right)$
@@
where $b(x)$ and $w_{b}$ are the base function (linear) and its base matrix. $\phi_{BS}(x)$ and $\phi_{RBF}(x)$ are B-spline and RBF functions, and $w_{s}$ is the common spline matrix for both functions.
Wav-KAN is a neural network architecture that integrates wavelet functions into Kolmogorov-Arnold Networks to address challenges in interpretability, training speed, robustness, and computational efficiency found in MLP and LiuKAN [7]. By efficiently capturing both high and low-frequency components of input data, Wav-KAN achieves a balance between accurately representing the data structure and avoiding overfitting. The authors used several wavelet types, including the DoG, Mexican hat, Morlet, and Shannon. In our paper, we use the DoG function to combine other functions to create function combinations. The formula for DoG is:
@@
For the simulation of function plots, we present them in Figure 2, where they are fitted to pass through four selected points. These plots provide a visual representation of the different function types and how they interact with the data points, helping to illustrate their shapes, which are not intended for pre-validating model performance trained on these functions. The functions analyzed include B-spline (3rd-degree), DoG, GRBF, RSWAF, and BSRBF (a combination of B-spline and GRBF). Among these, only the B-spline perfectly passes through all points, while DoG, BSRBF, and GRBF intersect with one point each, and RSWAF passes through none. It is important to note that a function passing through all specified points does not guarantee strong model performance in neural networks, as issues such as overfitting, poor generalization, and sensitivity to outliers may arise.